OpenStudy (anonymous):

. Which equation represents the line that passes through the points (6, –3) and (–4, –9)? A. y + 4 = 3 over 5 (x + 9) B. y + 4 = 5 over 3 (x + 9) C. y - 3 = 3 over 5 (x + 6) D. y + 3 = 3 over 5(x – 6)

OpenStudy (anonymous):

So first you have to find the slope: -9--3 over -4-6 So that's -6/-10 3/5

OpenStudy (anonymous):

Now for the equation... does it need to be in a certain form(like standard or point-slope)

OpenStudy (anonymous):

It has to look like one of these. A. y + 4 = 3 over 5 (x + 9) B. y + 4 = 5 over 3 (x + 9) C. y - 3 = 3 over 5 (x + 6) D. y + 3 = 3 over 5(x – 6)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that is point-slope

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

because: y--4=3/5(x--9) (negative and a negative = positive) y+4=3/5(x+9)

OpenStudy (anonymous):

Thanks you so much!

OpenStudy (anonymous):

No prob!

OpenStudy (anonymous):

Could you give me a medal?